OpenStudy (anonymous):

Find the equation of the tangent to the following parabola a) y^2=12x ; the tangent having slope -1/3

OpenStudy (solomonzelman):

The first thing you need to know is - where does the parabola have a slope -1/3, so that the tangent line would have the slope -1/3 ?

OpenStudy (solomonzelman):

\(y=\pm\sqrt{12x}\) \(y'=\pm\dfrac{6}{\sqrt{12x}}\) \(\dfrac{-1}{3}=\pm\dfrac{6}{\sqrt{12x}}\)

OpenStudy (solomonzelman):

we would need to take the bottom half of the parabola.

OpenStudy (anonymous):

where did u get that 6 frm \[\frac{ 6 }{ \sqrt{12x} }\]

OpenStudy (solomonzelman):

-1 / 3 = -6 / √(12x) 1 / 3 = 6 / √(12x) √(12x) = 18 12x = 324 x = 324/12 = 27

OpenStudy (solomonzelman):

y' = 1/ ( 2 √(12x)) * 12 = 6/ √(12)

OpenStudy (solomonzelman):

and ± comes from taking the square root on both sides

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

so I get that it is going through y = -(1/3) - 9 https://www.desmos.com/calculator/vpujilagiw

OpenStudy (anonymous):

so thats the equation?

OpenStudy (solomonzelman):

I forgot the x by the slope, oopsis

OpenStudy (solomonzelman):

-(1/3)x - 9